OpenStudy (marcelie):

Help Please !!1 A bacteria culture initially contains 20 cells and grows at a rate proportional to its size. After an hr the population has increased to 385. When will the population reach 20,000? Round ur answer a. t = 2.34 hrs b. t= 2.57 hrs c. t= 2.10 hrs d. t= 1.56 hrs e. 2.88 hrs

OpenStudy (marcelie):

@satellite73

OpenStudy (anonymous):

each hour it grows by a factor of \[\frac{385}{20}=\frac{77}{4}\]

OpenStudy (anonymous):

solve \[20\times \left(\frac{77}{4}\right)^t=20,000\]

OpenStudy (marcelie):